What are the dimensions of a triangle with a base of (n+10), height of 2n, and area of 75 units squared?
Answer by lwsshak3(11628)

What are the dimensions of a triangle with a base of (n+10), height of 2n, and area of 75 units squared?
***
1/2*base*height=area of a triangle
1/2(n+10)(2n)=75
n^2+10n=75
n^2+10n-75=0
(n+15)(n-5)=0
n=5
base of triangle=n+10=15 units
height of triangle=2n=10 units